# Resolving Interchange Reconnection Dynamics in a Fan-Spine-like Topology Observed by Solar Orbiter

### v1

@@ 3 Results
Figure 6 illustrates the spatial relationship between the fan-spine-like structure and the surrounding magnetic field. To enhance the visibility of this structure, HRIEUV 174 Å images are superimposed on a sequence of FSI 174 Å images, see Figure 6 (b1)(b3). Additionally, running difference images from the FSI 304 Å sequence, shown at 10-minute intervals, depict the process of plasma outflow. To quantify the plasma motion, two cuts were placed, as shown in Figure 6 (a). It should be noted that Slice 1 is positioned along the path where the plasma ejection reaches its maximum height within the fan-spine-like structure, while Slice 2 is established in the direction of the eruption of filament 1 along the side of this structure. By calculating the light curve within the yellow dashed box in Figure 6 (a), we found that the peak of this intensity curve corresponds to the eruption of filament 1. Based on the time-distance diagram of Figure 6 (e), we determined that the time and speed of the erupted filament 1 are about 22:18 UT and 186 km s-1, respectively. The average speed of the recurring plasma ejections is 64 km s-1. The highest plasma ejection occurred around 22:23 UT (see Figure 6 (d)), corresponding to the eruption of filament 2 in Figures 3 (d)(f) and (i3). In the FSI 304 Å running difference images, it can be observed that the plasma material flows along the closed field lines and finally impacts the solar surface, as denoted by white arrows in Figure 6 (c2) and (c3). No disturbance signals were observed in the HRIEUV 174 Å or FSI 174 Å images, indicating that the temperature of the material flows is not high relative to the coronal loop. Due to the low temporal resolution of the FSI, it is not possible to distinguish which filament eruption the material propagating along the closed magnetic field lines in Figure (c2)(c3) originates from. However, these results indicate that the reconnection occurring in the low corona can contribute to supplying material and energy to the high corona. Additionally, the ejected plasma flows subsequently fall down, possibly corresponding to coronal rain within the coronal loops. The ongoing coronal rain can indicate the spatial distribution of heating of the coronal loops (Antolin et al., 2010; Antolin & Froment, 2022).
@@
In Figure 7, to investigate the periodic oscillation of plasma outflows along the jet spire and the curtain feature, we applied wavelet and Fourier analysis to four light curves, following the methodology introduced by Torrence & Compo (1998)and Auchère et al. (2016). The positions from which the light curves were extracted are marked by the red dashed lines in Figure 3 (i3) and Figures 4 (i1)(i3). We chose to fit the fast Fourier transform (FFT) of each light curve using the following power-law equation:
@@
where is the Fourier frequency; A, B, and C are constants; the second term accounts for the possible presence of pulses in the time series; is the Kappa function:
@@
where is the width of the power spectral distribution (PSD) and defines the extent of its high-frequency wing.
@@
As shown in Figures 7 (a3)(d3), the global 99% Fourier confidence level exhibits higher values compared to the local and global 99% wavelet confidence levels. The global 99% wavelet confidence levels are higher than the local 99% wavelet confidence level. This result is consistent with findings reported in previous studies (e.g., Auchère et al., 2016; Kayshap et al., 2020; Li, 2022; Mandal et al., 2022). In Figure 7 (a3), (c3), and (d3), a slope variation of the Fourier spectrum forms a peak around , corresponding to a period of 200 s. This indicates that periodic interchange reconnection occurs on the curtain structure (Figure 7 (b3), (c3), and (d3)), comparable to the periodicity of interchange reconnection at the spire of the jet (Figure 7 (a3)), with both being approximately 200 s. It should be noted that during the eruption, the upward movement of cold plasma obscures the hot plasma outflows. This may explain the lack of a distinct periodicity in Figure 7 (b3). In this scenario, the periodicity of the interchange reconnection may also be underestimated.
@@ 4 Conclusion and Discussion
Using high spatial resolution observations from the Extreme Ultraviolet Imager (EUI) onboard the Solar Orbiter, we report direct imaging of three current sheets (CSs), associated with recurrent interchange reconnection near the null point within a fan-spine-like system. Three current sheets occurring near the null point are modulated by emerged small magnetic structures. CS1 is generated by the interaction between emerging magnetic loops and the ambient magnetic field. The X point at the reconnection site and plasma blobs within CS1 are observed. CS2 appears to be influenced by the eruption of filament 1 on the side of the fan-spine-like topology, lasting about 6 minutes with a maximum length of 5 Mm and a width of 0.8 Mm. The formation mechanism of CS3 is similar to the flare current sheet in the breakout jet model. CS3 lasts 17 minutes with a maximum length of 7.4 Mm and a width of 0.6 Mm. The estimated reconnection rates are approximately 0.12 to 0.2 for the CS2 and 0.12 to 0.17 for the CS3. A curtain-like feature with a width of 1.7 Mm is presented during the dynamic evolution of the fan-spine system. Periodic interchange reconnection occurs in the curtain, with a periodicity comparable to that of interchange reconnection at the jet spire, both being about 200 s. Additionally, the plasma outflows from periodic reconnections, regardless of where they occur, nearly encompass the evolution of the fan-spine system, exhibiting a gradual physical process similar to that of interchange reconnection in a pseudo-streamer. We suggest that the dynamics of recurrent magnetic reconnection in such a small fan-spine system might be more complex than we recognized before. Our results reveal a self-similarity between small-scale fan-spine structures and large-scale pseudo-streamer structures.
@@
In our research, we found that during the dynamic evolution of the fan-spine system, complex magnetic field interchange reconnection occurs near the null point. Specifically, the appearance of CS1 results from the interchange reconnection between the newly emerging magnetic flux and background fields (Figure 2 ); the transition from CS2 to CS3 is accompanied by exchanges in reconnection outflows and inflows (Figure 3); interchange reconnection occurs between the closed magnetic loop under the fan dome and the surrounding open field, forming a curtain-like structure in the reconnection outflow region of CS3 (Figure 4); and periodic reconnection occur near the null point, including on the jet spire and the curtain feature. These results suggest that the magnetic reconnection in a fan-spine topology seems to be more complex than single null-point reconnection, possibly involving QSL reconnection or separators (i.e., a series of nulls). As shown in Duan et al. (2024b), they found that two null points can be connected by an elongated sheetlike quasi-separatrix layer (QSL) through the magnetic null point tracing method and the calculation of the 3D squashing factor Q. This result indicates the possible existence of multiple magnetic null points in the fan-spine system.
@@
High-resolution EUI 174 Å images reveal that persistent plasma outflows occur near the magnetic null point. The velocity of these plasma outflows ( 64-90 km s-1) is relatively lower compared to the filament 1 eruption (i.e., 186 km s-1), indicating that gentle reconnection occurs persistently near the magnetic null point. The recurrent reconnection that persistently occurs near the magnetic null point in the fan-spine configuration, resulting in plasma outflows, has also been reported by previous studies (Cheng et al., 2023; Duan et al., 2024b; Kumar et al., 2024). This process may resemble the simulations by Nóbrega-Siverio & Moreno-Insertis (2022); Nóbrega-Siverio et al. (2023), where stochastic granular motions are sufficient to stress the small-scale fan-spine structure, triggering sustained reconnection at the coronal null point, which naturally results in persistent jetting activity manifested in the form of recurrent, gentle plasma outflows. The gentle reconnection or gentle plasma outflows most likely result from the interchange reconnection between magnetic flux under the fan dome and the ambient open magnetic field lines. Such gradual physical processes are similar to the interchange reconnection occurring in the pseudo-streamer structure (e.g., Masson et al., 2014) and may play a critical role in the continuous energy and material injection into the high corona.
@@
The observation that sparked our interest in this event was the identification of two distinct current sheets (CS2 and CS3) at disparate locations. It is pertinent to note that changes in magnetic pressure or gas pressure may lead to reconnection reversals (Murray et al., 2009), resulting in the formation of oscillatory current sheets around the null point, as demonstrated by Hong et al. (2019). Their study focused on oscillatory magnetic reconnection preceding a jet within the breakout current sheet. Here, the appearance of CS2 and CS3 in sequence also indicates the reconnection reversal. However, it should be attributed to the dynamics described by the breakout jet model, different from the oscillatory reconnection. In simulations conducted by Wyper et al. (2017), shear facilitated the accumulation of free energy within the filament channels along the polarity inversion line of the fan-spine magnetic configuration. As shear progressed, the overlying magnetic field expanded upward, leading to the formation of a breakout current sheet (BCS) at the null point. In this event, the appearance of CS2 is influenced by the eruption of filament 1 on the side of the fan-spine-like structure. Reconnection at the CS2 then eliminates the restraining fields, resulting in the eruptions of another filament (filament 2) (e.g., Wyper et al., 2017, 2018; Kumar et al., 2018; Sun et al., 2023; Duan et al., 2024a, b; Yang et al., 2024b). Filament 2 then rapidly reconnected with the external open field, inducing explosive interchange reconnection within the CS3 left in its wake, thereby producing the post-flare loops. The transition from CS2 to CS3 in this event is similar to the breakout jet model (Wyper et al., 2017). In the simulation studies (e.g., Scott et al., 2021), the reversal or transition of such current sheets seems not to have been considered. Our observations may also indicate that the dynamics of large-scale pseudo-streamer structures could undergo similar complex processes, which might be a possible reason for the appearance of complex structures in the solar wind.
@@
It should be noted that our calculations of the width of the current sheet may overestimate the actual width because the features we observed are essentially plasma sheets embedded within the current sheet (e.g., Lin et al., 2015). What we are calculating here is the upper limit of the width of the current sheet. Considering that the thermal halo effect has minimal impact, the lengths of the current sheets are generally comparable to the lengths of the plasma sheets. Therefore, the calculated value of reconnection rate is likely overestimated, representing only an upper limit of the true reconnection rate. However, the overall trend reflects the evolutionary characteristics of the reconnection rate.
@@
In the large-scale pseudo-streamer structures, the separatrix curtain formed by the separatrix fan surface from the central null is orthogonal to the other two surfaces and is bounded by their respective spine field lines (Titov et al., 2011). Compared to the fan-spine topology (often described as a separatrix dome with a single null point), separatrix curtains involve multiple null points connected by separators.In fact, considering the three-dimensional invariance of the fan-spine structure in a two-dimensional geometry, the three-dimensional null point is replaced with a separator, and the spine line is replaced with a spine surface. Within the three-dimensional magnetic reconnection framework, reconnection is not confined to null points but occurs throughout a finite diffusion region (Pontin & Priest, 2022). In this scenario, the fan-spine structure resembles a scaled-down version of the pseudostreamer structure (e.g., Lynch & Edmondson, 2013; Wyper et al., 2024). In the small-scale fan-spine topology, the presence of multiple nulls and connecting separators allows for more frequent and complex interchange reconnection between open and closed magnetic fluxes, similar to the pseudo-streamer simulated by Scott et al. (2021). Consistent with our previous findings (Duan et al., 2024b), our current observations reveal the occurrence of complex and frequent interchange reconnection in the coronal jets between open and closed magnetic fluxes, likely mediated by multiple null points. Concurrently, the coronal jet expelled plasma with a curtain-like jet spire and complex current sheet evolution. These results provide evidence that this coronal jet occurs in a small-scale sepatatrix curtain, resembling a scaled-down version of the pseudostreamer structure. Similarly, Mandal et al. (2022) recently reported analogous phenomena, including a curtain-like jet spire and numerous complex morphological changes, in a highly dynamic, small-scale event observed by EUI in a polar coronal hole. Moreover, Li et al. (2025a) also reconstructed a separatrix curtain-like magnetic topology in small-scale recurrent jets. These high-resolution EUI imaging observations jointly reinforce the view of the across-scale self-similarity of magnetic separatrix curtains and their associated interchange reconnection processes from the large-scale pseudo-streamers down to the tiny coronal jets.

### v2

@@ 3 Results
Figure 6 illustrates the spatial relation between the fan-spine-like structure and the surrounding magnetic field. For clarity, we superposed HRIEUV 174 Å images on a sequence of FSI 174 Å images (see Figs. 6(b1)6(b3)). Additionally, the running-difference images from the FSI 304 Å sequence, shown at 10-minute intervals, show the process of plasma outflow. To quantify the plasma motion, two cuts were placed, as shown in Fig. 6(a). Slice 1 was positioned along the path where the plasma ejection reached its maximum height within the fan-spine-like structure, and slice 2 was established in the direction of the eruption of filament 1 along the side of this structure. By calculating the light curve within the dashed yellow box in Fig. 6(a), we found that the peak of this intensity curve corresponds to the eruption of filament 1. Based on the time-distance diagram of Fig. 6(e), we determined that the time and speed of the erupted filament 1 are about 22:18 UT and 186 km s-1. The average speed of the recurring plasma ejections is 64 km s-1. The highest plasma ejection occurred around 22:23 UT (see Fig. 6(d)), corresponding to the eruption of filament 2 in Figs. 3(d)3(f) and (i3). The FSI 304 Å running-difference images show that the plasma material flows along the closed field lines and finally impacts the solar surface, as denoted by white arrows in Figs. 6(c2) and 6(c3). No disturbance signals were observed in the HRIEUV 174 Å or FSI 174 Å images, indicating that the temperature of the material flows was not high relative to the coronal loop. The low temporal resolution of the FSI prevents us from distinguishing which filament eruption the material that propagated along the closed magnetic field lines in Figs. (c2)(c3) originated. These results indicate, however, that the reconnection in the low corona can supply material and energy to the high corona. Additionally, the ejected plasma flows subsequently fall down, which might correspond to coronal rain within the coronal loops. The ongoing coronal rain indicates the spatial distribution of the coronal-loop heating (Antolin et al., 2010; Antolin & Froment, 2022).
@@
To investigate the periodic oscillation of plasma outflows along the jet spire and the curtain feature, we applied a wavelet and Fourier analysis in Fig. 7 to four light curves, following the method introduced by Torrence & Compo (1998)and Auchère et al. (2016). The positions from which the light curves were extracted are marked by the dashed red lines in Fig. 3(i3) and Figs. 4(i1)(i3). We chose to fit the fast Fourier transform (FFT) of each light curve using the following power-law equation:
@@
where is the Fourier frequency, A, B, and C are constants, the second term accounts for the possible presence of pulses in the time series, and is the Kappa function,
@@
where is the width of the power spectral distribution (PSD), and defines the extent of its high-frequency wing.
@@
Figs. 7(a3)7(d3) show that the global 99% Fourier confidence level is higher than the local and global 99% wavelet confidence levels. The global 99% wavelet confidence levels are higher than the local 99% wavelet confidence level. This result is consistent with findings reported in previous studies (e.g., Auchère et al., 2016; Kayshap et al., 2020; Li, 2022; Mandal et al., 2022). In Figs. 7(a3), 7(c3), and 7(d3), a slope variation in the Fourier spectrum forms a peak around , corresponding to a period of 200 s. This indicates that periodic interchange reconnection occurred in the curtain structure (Figs. 7(b3), 7(c3), and 7(d3)), which is comparable to the periodicity of the interchange reconnection at the jet spire (Fig. 7(a3)), with both being approximately 200 s. During the eruption, the upward movement of cold plasma obscures the hot plasma outflows. This might explain the lack of a distinct periodicity in Fig. 7(b3). In this scenario, the periodicity of the interchange reconnection is maybe also underestimated.
@@ 4 Conclusion and discussion
Using high spatial resolution observations from the Extreme Ultraviolet Imager (EUI) on board the Solar Orbiter, we reported the direct imaging of three current sheets (CSs) associated with recurrent interchange reconnection near the null point within a fan-spine-like system. The three current sheets occurring near the null point were modulated by emerged small magnetic structures. CS1 was generated by the interaction between emerging magnetic loops and the ambient magnetic field. An X point at the reconnection site and plasma blobs within CS1 were observed. CS2 appeared to be affected by the eruption of filament 1 on the side of the fan-spine-like topology, lasting about 6 minutes with a maximum length of 5 Mm and a width of 0.8 Mm. The formation mechanism of CS3 is similar to the flare current sheet in the breakout jet model. CS3 lasted 17 minutes, with a maximum length of 7.4 Mm and a width of 0.6 Mm. The estimated reconnection rates are approximately 0.12 to 0.2 for the CS2 and 0.12 to 0.17 for the CS3. A curtain-like feature with a width of 1.7 Mm arose during the dynamic evolution of the fan-spine system. Periodic interchange reconnection occurred in the curtain, with a periodicity comparable to that of interchange reconnection at the jet spire, both being about 200 s. Additionally, the plasma outflows from periodic reconnections, regardless of where they occurred, nearly encompassed the evolution of the fan-spine system and showed a gradual physical process similar to that of interchange reconnection in a pseudostreamer. We suggest that the dynamics of recurrent magnetic reconnection in such a small fan-spine system might be more complex than we realized before. Our results revealed a self-similarity between small-scale fan-spine structures and large-scale pseudostreamer structures.
@@
We found that during the dynamic evolution of the fan-spine system, complex magnetic field interchange reconnection occurred near the null point. Specifically, the appearance of CS1 resulted from the interchange reconnection between the newly emerging magnetic flux and background fields (Fig. 2); the transition from CS2 to CS3 was accompanied by exchanges in reconnection outflows and inflows (Fig. 3); interchange reconnection occurred between the closed magnetic loop under the fan dome and the surrounding open field, forming a curtain-like structure in the reconnection outflow region of CS3 (Fig. 4); and periodic reconnection occurred near the null point, including on the jet spire and the curtain feature. These results suggest that the magnetic reconnection in a fan-spine topology seems to be more complex than single null-point reconnection, possibly involving QSL reconnection or separators (i.e., a series of nulls). Duan et al. (2024b) reported that two null points can be connected by an elongated sheetlike quasi-separatrix layer (QSL) through the magnetic null-point-tracing method and the calculation of the 3D squashing factor Q. This result indicates the possible existence of multiple magnetic null points in the fan-spine system.
@@
High-resolution EUI 174 Å images revealed that persistent plasma outflows occur near the magnetic null point. The velocity of these plasma outflows ( 64-90 km s-1) is somewhat lower than that of the filament 1 eruption (i.e., 186 km s-1), indicating that gentle reconnection occurs persistently near the magnetic null point. Recurrent reconnection that persistently occurs near the magnetic null point in the fan-spine configuration, resulting in plasma outflows, was also reported in previous studies (Cheng et al., 2023; Duan et al., 2024b; Kumar et al., 2024). This process resembles the simulations by Nóbrega-Siverio & Moreno-Insertis (2022); Nóbrega-Siverio et al. (2023), where stochastic granular motions were sufficient to stress the small-scale fan-spine structure, triggering sustained reconnection at the coronal null point, which naturally resulted in persistent jetting activity in the form of recurrent gentle plasma outflows. Gentle reconnection or gentle plasma outflows most likely result from the interchange reconnection between magnetic flux under the fan dome and the ambient open magnetic field lines. These gradual physical processes are similar to the interchange reconnection that occurs in the pseudostreamer structure (e.g., Masson et al., 2014) and probably plays a critical role in the continuous energy and material injection into the high corona.
@@
The observation that sparked our interest in this event was the identification of two distinct current sheets (CS2 and CS3) at disparate locations. It is pertinent to note that changes in magnetic pressure or gas pressure can lead to reconnection reversals (Murray et al., 2009), resulting in the formation of oscillatory current sheets around the null point, as demonstrated by Hong et al. (2019). Their study focused on oscillatory magnetic reconnection preceding a jet within the breakout current sheet. Here, the appearance of CS2 and CS3 in sequence also indicates the reconnection reversal. This should be attributed to the dynamics described by the breakout jet model, however, which is different from the oscillatory reconnection. In simulations conducted by Wyper et al. (2017), shear facilitated the accumulation of free energy within the filament channels along the polarity inversion line of the fan-spine magnetic configuration. As the shear progressed, the overlying magnetic field expanded upward and formed a breakout current sheet (BCS) at the null point. In this event, the appearance of CS2 was affected by the eruption of filament 1 on the side of the fan-spine-like structure. Reconnection at the CS2 then eliminated the restraining fields, resulting in eruptions of another filament (filament 2; e.g., Wyper et al., 2017, 2018; Kumar et al., 2018; Sun et al., 2023; Duan et al., 2024a, b; Yang et al., 2024b). Filament 2 then rapidly reconnected with the external open field, inducing explosive interchange reconnection within the CS3 left in its wake, thereby producing the post-flare loops. The transition from CS2 to CS3 in this event is similar to the breakout jet model (Wyper et al., 2017). In the simulation studies (e.g., Scott et al., 2021), the reversal or transition of these current sheets was apparently not considered. Our observations also indicate that the dynamics of large-scale pseudostreamer structures might undergo similar complex processes, which might be a possible reason for the appearance of complex structures in the solar wind.
@@
Our calculations of the width of the current sheet might overestimate the actual width because the features we observed were essentially plasma sheets embedded within the current sheet (e.g., Lin et al., 2015). When we consider that the thermal halo effect has very little impact, the lengths of the current sheets are generally comparable to the lengths of the plasma sheets. The calculated reconnection rate is therefore likely overestimated and only represents an upper limit of the true reconnection rate. The overall trend reflects the evolutionary characteristics of the reconnection rate, however.
@@
In large-scale pseudostreamer structures, the separatrix curtain formed by the separatrix fan surface from the central null is orthogonal to the other two surfaces and is bounded by their respective spine field lines (Titov et al., 2011). Compared to the fan-spine topology (often described as a separatrix dome with a single null point), separatrix curtains involve multiple null points that are connected by separators. When we consider the three-dimensional invariance of the fan-spine structure in a two-dimensional geometry, the three-dimensional null point is replaced with a separator, and the spine line is replaced with a spine surface. Within the three-dimensional magnetic reconnection framework, reconnection is not confined to null points, but occurs throughout a finite diffusion region (Pontin & Priest, 2022). In this scenario, the fan-spine structure resembles a scaled-down version of the pseudostreamer structure (e.g., Lynch & Edmondson, 2013; Wyper et al., 2024). In the small-scale fan-spine topology, multiple nulls and connecting separators allow a more frequent and complex interchange reconnection between open and closed magnetic fluxes, similar to the pseudostreamer simulated by Scott et al. (2021). Consistent with our previous findings (Duan et al., 2024b), our current observations revealed the occurrence of a complex and frequent interchange reconnection in the coronal jets between open and closed magnetic fluxes, likely mediated by multiple null points. Concurrently, the coronal jet expelled plasma with a curtain-like jet spire and a complex current sheet evolution. These results provide evidence that this coronal jet occurred in a small-scale separatrix curtain, resembling a scaled-down version of the pseudostreamer structure. Similarly, Mandal et al. (2022) recently reported analogous phenomena, including a curtain-like jet spire and numerous complex morphological changes, in a highly dynamic small-scale event observed by EUI in a polar coronal hole. Moreover, Li et al. (2025a) also reconstructed a separatrix curtain-like magnetic topology in small-scale recurrent jets. These high-resolution EUI imaging observations jointly reinforce the view of the across-scale self-similarity of magnetic separatrix curtains and their associated interchange reconnection processes from the large-scale pseudostreamers down to tiny coronal jets.